A table and two chairs together cost $400. If by selling the chairs at 10% loss and the table at 10% profit, a total of 5% profit is made. What is the cost price of a chair?


Answer by ankor@dixie-net.com(22740)

A table and two chairs together cost $400.
If by selling the chairs at 10% loss and the table at 10% profit, a total of 5% profit is made.
What is the cost price of a chair?
:
Let c = cost of 1 chair
then
2c = cost of 2 chairs
and
(400-2c) = cost of the table
:
Find 5% of $400 = $20, the profit
:
Table profit - chair loss = $20
.10(400-2c) - .10(2c) = 20
40 - .2c - .2c = 20
-.4c = 20 - 40
-.4c = -20
c = -20/-.4
c = + $50 for each chair
:
:
See if that checks out
Table cost 400 - 2(50) = $300
:
Table profit: .10(300) = 30
Chair loss: -.10(2*50) = -10
----------------------------
The total profit = $20 which is 5% of $400